# Exceptional Tannaka groups only arise from cubic threefolds

### v2

@@ 8.1. Main result
We will rule out $E_{7}$ via a general criterion for a subvariety to have a big Tannaka group if the alternating or symmetric convolution square has few summands. We go back to the notation introduced in section 1.1, so let $A$ be an abelian variety of dimension $g$ over an algebraically closed field $k$ of characteristic zero, and consider a smooth integral subvariety $X\subset A$ of dimension $d$. We again fix a fiber functor $\omega\colon\langle\delta_{X}\rangle\to\operatorname{Vect}(\mathbb{F})$ and denote the corresponding Tannaka group by
@@
where the involution of $\delta_{X}*\delta_{X}$ permuting the two factors acts as $\operatorname{id}$ on $\operatorname{Sym}^{2}(\delta_{X})$ and $-\operatorname{id}$ on $\operatorname{Alt}^{2}(\delta_{X})$. Define
@@
If $X\subset A$ is symmetric up to translation in the sense that $X=a-X$ for some point $a\in A(k)$, then $\delta_{a}\subset T_{+}(\delta_{X})$ enters as a direct summand. When $X$ is nondivisible the point $a$ is unique and we define
@@
for a unique complex $S_{+}(\delta_{X})$. The skyscraper summand defines a bilinear form $\theta\colon V\otimes V\to\omega(\delta_{a})$ which is symmetric if $d$ is even and alternating if $d$ is odd. Hence the derived subgroup $G_{X,\omega}^{\ast}$ of the connected component of $G_{X,\omega}$ is contained in $\operatorname{SO}(V,\theta)$ is $d$ is even and in $\operatorname{Sp}(V,\theta)$ if $d$ is odd. To have a uniform notation, when $X$ is not symmetric up to translation we set
@@
$G_{X,\omega}^{\ast}=\begin{cases}\operatorname{SL}(V)&\text{if $X$ is not symmetric up to translation},\\ \operatorname{SO}(V,\theta)&\text{if $X$ is symmetric up to translation and $d$ is even},\\ \operatorname{Sp}(V,\theta)&\text{if $X$ is symmetric up to translation and $d$ is odd}.\end{cases}$
Note that in comparison to Larsen’s alternative in [JKLM23, lemma 3.7], we here only need an assumption on the symmetric or alternating square rather than on the full tensor square. The rest of this section will be devoted to the proof of the above criterion, see section 8.5. But before coming to the proof, let us show how theorem 8.1 rules out the exceptional Tannaka group $E_{7}$:
@@
Suppose $G_{X,\omega}^{\ast}\simeq E_{7}$. If $X\subset A$ descends under an isogeny $A\to B$, then its image under this isogeny is still smooth nondegenerate and irreducible of the same dimension with Tannaka group $E_{7}$. Hence in what follows we assume that $X$ is nondivisible. Then by [Krä22, cor. 1.10] the representation $V=\omega(\delta_{X})$ of $G_{X,\omega}^{\ast}$ is minuscule and hence isomorphic to the $56$-dimensional irreducible representation of $E_{7}$. By the highest weight theory of the exceptional group $E_{7}$ then
@@
where $W$ is irreducible and $\mathbf{1}$ is the trivial representation of dimension one. In particular, there is an alternating bilinear form on $V$ which is invariant under the action $E_{7}$. It follows that $d$ is odd and, after a suitable translation, that $X$ is symmetric. Then
@@
and $S_{+}(\delta_{X})$ is perverse by the first part of theorem 8.1 and it corresponds to the representation $\omega(S_{+}(\delta_{X}))\simeq W$. The latter is irreducible, so $S_{+}(\delta_{X})$ is a simple perverse sheaf up to negligible direct summands. The second part of theorem 8.1 then says that $G_{X,\omega}^{\ast}$ is the full symplectic group $\operatorname{Sp}(V,\theta)$, a contradiction. ∎
@@ 8.2. Topological preliminaries
In this section we work over $k=\mathbb{C}$. For the proof of theorem 8.1 we will detect perverse direct summands of a convolution product from the top cohomology of the fibers of the sum morphism, using the following well-known fact:
@@
The above lemma is about the cohomology with constant coefficients, so we want a criterion to decide when the perverse intersection complex on a singular variety is the constant sheaf. A complex variety $W$ is called a rational homology manifold if
@@ 8.3. Kashiwara’s estimate
In this section we continue to work over $k=\mathbb{C}$. We will control negligible summands in convolution products in terms of characteristic varieties. As in [JKLM23, def. 2.2], we attach to any subvariety $Y\subset A$ its projective conormal variety
@@
the union of the characteristic varieties of all its perverse cohomology sheaves. We can then control the characteristic varieties of convolution products $K=P*P$ by the following incarnation of Kashiwara’s estimate:
@@
Let $P_{1},P_{2}\in\textup{Perv}(A)$. If the Gauss maps $\mathbb{P}\operatorname{Char}(P_{i})\to\mathbb{P}_{A}$ are both finite, then for any perverse direct summand
$Q\;\subset\;^{p}\mathcal{H}^{i}(P_{1}*P_{2})$
with $i\in\mathbb{Z}$, the support $Y=\operatorname{Supp}(Q)$ is either of general type or $Y=A$.
@@ 8.4. Decomposition of the convolution square
The main point in the proof of theorem 8.1 will be to deduce from our assumption about the symmetric or alternating square the decomposition of the full tensor square. To explain this we go back to the notation of section 8.1. Let $T_{-}(\delta_{X})$ denote the direct summand complementary to $T_{+}(\delta_{X})$ in the decomposition of the convolution square $\delta_{X}\ast\delta_{X}$, so that
@@
(2) Part (1) implies that $S_{+}$ is perverse. To understand this more geometrically, consider the factorization
${X\times X}$${X+X}$${\operatorname{Sym}^{2}X}$$\scriptstyle{a}$$\scriptstyle{\pi}$$\scriptstyle{f}$
where $\pi$ denotes the quotient map. Write
@@
and by lemma 8.4 this sheaf is the perverse intersection complex on $\operatorname{Sym}^{2}X$ up to a shift. The definition of the commutativity constraint for the convolution product moreover shows
@@
The decomposition theorem together with the fact that $f$ is generically finite implies that $\delta_{X+X}\subset T_{+}(\delta_{X})$. Since by assumption $S_{+}(\delta_{X})$ is a simple perverse sheaf, we see that
@@
Then $f_{*}L_{+}\simeq\mathcal{H}^{-2d}(T_{+}(\delta_{X}))$ has generic rank one on its support $X+X$, which means that
$f\colon\quad\operatorname{Sym}^{2}X\;\longrightarrow\;X+X$
is birational. But this birationality implies that also $\mathcal{H}^{-2d}(T_{-}(\delta_{X}))\simeq f_{*}L_{-}$ has generic rank one on its support $X+X$. The structure of perverse sheaves forces
@@
where $Q$ is a simple perverse sheaf of generic rank one on its support $X+X$ and $R$ is either zero or a semisimple perverse sheaf with strictly smaller support. Note that $R$ cannot be a skyscraper sheaf, as $\delta_{X}\ast\delta_{X}$ contains by adjunction only skyscrapers of the form $\delta_{a}$ for $a\in A(\mathbb{C})$ such that $X=p-X$. Since $X$ is nondivisible, there is at most one such skyscraper sheaf and it enters $T_{+}(\delta_{X})$.
@@
by lemma 8.3. It follows that we have
@@
8.5. Proof of theorem 8.1
Part (1) of proposition 8.7 says that $S_{+}(V)$ is a perverse sheaf without negligible direct summands, so it only remains to show that if $S_{+}(\delta_{X})$ is a simple perverse sheaf, then
$G_{X,\omega}^{\ast}=\begin{cases}\operatorname{SL}(V)&\text{if $X$ is not symmetric up to translation},\\ \operatorname{SO}(V,\theta)&\text{if $X$ is symmetric up to translation and $d$ is even},\\ \operatorname{Sp}(V,\theta)&\text{if $X$ is symmetric up to translation and $d$ is odd}.\end{cases}$
But this follows from part (2b) of proposition 8.7 and from the version of Larsen’s alternative given in [JKLM23, lemma 3.7]. For convenience we recall the simple argument when the subvariety $X\subset A$ is symmetric up to translation, which is the only one needed for the application to $E_{7}$. The symmetry of $X$ implies that the representation $V$ is self-dual and we have an isomorphism $V\otimes V\simeq\operatorname{End}(V)$. Let $\mathfrak{h}\subset\operatorname{End}(V)$ be the Lie algebra of $\operatorname{SO}(V,\theta)$ if $d$ is even and of $\operatorname{Sp}(V,\theta)$ if $d$ is odd. Via the preceding isomorphism we have
@@
The representation $\mathfrak{h}$ of $G_{X,\omega}$ therefore corresponds via Tannaka duality to the perverse sheaf $S_{-}(\delta_{X})$, which is simple by part (2b) of proposition 8.7. Thus $\mathfrak{h}$ is an irreducible representation of the group $G_{X,\omega}$. On the other hand, the adjoint representation $\mathfrak{g}$ of this group is a subrepresentation of $\mathfrak{h}$, thus $\mathfrak{g}=\mathfrak{h}$. ∎
@@ Appendix A Hodge number estimates
$(-1)^{d-p}\chi(X,\Omega^{p}_{X})\;\geqslant\;\begin{cases}\dim V-d+1&\text{for\leavevmode\nobreak\ $p\,\in\,\{0,d\}$},\\[1.00006pt] 2&\text{for\leavevmode\nobreak\ $p\,\in\,\{1,d-1\}$},\\[1.00006pt] 1&\text{for\leavevmode\nobreak\ $2\leqslant p\leqslant d-2$}.\end{cases}$
@@
When $p=d-1$, let $G:=\operatorname{Ker}(\delta_{1})$ so the complex $\mathbb{L}^{p}_{X,V}$ becomes an exact sequence
$00\;\longrightarrow\;G\;\longrightarrow\;\mathcal{O}_{\mathbb{P}}(-d+1)\otimes\textup{H}^{1}(X,\Omega^{d-1}_{X})\;\longrightarrow\;\cdots\longrightarrow\;\mathcal{O}_{\mathbb{P}}\otimes\textup{H}^{d}(X,\Omega^{d-1}_{X})\;\longrightarrow\;$ 0 .
@@
$0\longrightarrow\mathcal{O}_{\mathbb{P}}(-d)\otimes\textup{H}^{0}(X,\Omega^{1}_{X})\longrightarrow G\longrightarrow\mathcal{O}_{\mathbb{P}}(j)\longrightarrow 0$
@@
$-\chi(X,\Omega^{d-1}_{X})\;=\;\operatorname{rk}G\;\geqslant 2.$
@@
$\textup{H}^{i}(\mathbb{P},G^{\vee}(j))\;=\;0\quad\text{for all\leavevmode\nobreak\ $j\in\mathbb{Z}$ and all\leavevmode\nobreak\ $i\in\{1,2,\dots,\dim V-d-1\}$}.$
If $\operatorname{rk}G\leqslant\dim V-d$, then by the Evans-Griffiths theorem [Laz04, ex. 7.3.10] it would follow that $G$ is a direct sum of line bundles. This is seen to be impossible by applying [Lom13, lemma 2.2 (ii)] with $e=d$, $q=\dim V$ and $a=-j$. Therefore, we have
$\operatorname{rk}G\;\geqslant\;\dim V-d+1$
and the claim follows since from the above exact sequence $\operatorname{rk}G=\chi(X,\Omega_{X}^{d})$. ∎
@@
$(-1)^{d-p}\chi(X,\Omega^{p}_{X})\;\geqslant\;\begin{cases}g-d+1&\text{for\leavevmode\nobreak\ $p\,\in\,\{0,d\}$},\\[1.00006pt] 2&\text{for\leavevmode\nobreak\ $p\,\in\,\{1,d-1\}$},\\[1.00006pt] 1&\text{for\leavevmode\nobreak\ $2\leqslant p\leqslant d-2$}.\end{cases}$
@@
Since $X\subset A$ has ample normal bundle, we know by [Laz04, prop. 6.3.10(i)] that for every
$0\;\neq\;\omega\;\in\;V\;:=\;\textup{H}^{0}(A,\Omega^{1}_{A})\;\subset\;\textup{H}^{0}(X,\Omega^{1}_{X}),$
the zero locus $Z(\omega)\subset X$ is finite. Hence, $m(V)=d$, and theorem A.1 applies. ∎

### v3

@@ 8.1. Main result
We will rule out $E_{7}$ via a general criterion for a subvariety to have a big Tannaka group if the alternating or symmetric convolution square has few summands. We go back to the notation introduced in section˜1.1, so let $A$ be an abelian variety of dimension $g$ over an algebraically closed field $k$ of characteristic zero, and consider a smooth integral subvariety $X\subset A$ of dimension $d$. We again fix a fiber functor $\omega\colon\langle\delta_{X}\rangle\to\operatorname{Vect}(\mathbb{F})$ and denote the corresponding Tannaka group by
@@
where the involution of $\delta_{X}*\delta_{X}$ permuting the two factors acts as $\operatorname{id}$ on $\operatorname{Sym}^{2}(\delta_{X})$ and $\operatorname{id}$ on $\operatorname{Alt}^{2}(\delta_{X})$. Define
@@
If $X\subset A$ is symmetric up to translation in the sense that X=aXX=a-X for some point $a\in A(k)$, then the fiber of the sum morphism $\sigma\colon X\times X\to A$ over this point is the antidiagonal $\sigma^{-1}(a)=\{(x,a-x)\mid x\in X\}$. By proper base change and the decomposition theorem [BBDG18, th. 6.2.5], we then have a skyscraper direct summand $\delta_{a}\subset\delta_{X}*\delta_{X}$, and a closer look at the parity shows that $\delta_{a}\subset T_{+}(\delta_{X})$; see [JKLM25, sect. 1.2] [KW15a, lemma 2.1]. When $X$ is nondivisible, the point $a$ is unique and we define
@@
for a unique complex $S_{+}(\delta_{X})$. By the dimension formula (2.1), the skyscraper direct summand corresponds to a one-dimensional representation of the Tannaka group, and the projection onto it defines a bilinear form $\theta\colon V\otimes V\to\omega(\delta_{a})$ which is symmetric if $d$ is even and alternating if $d$ is odd. Hence the derived subgroup $G_{X,\omega}^{\ast}$ of the connected component of $G_{X,\omega}$ is contained in $\operatorname{SO}(V,\theta)$ is $d$ is even and in $\operatorname{Sp}(V,\theta)$ if $d$ is odd. To have a uniform notation, when $X$ is not symmetric up to translation we set
@@
[…]
Note that in comparison to Larsen’s alternative in [JKLM25, lemma 3.7], we here only need an assumption on the symmetric or alternating square rather than on the full tensor square. The rest of this section will be devoted to the proof of the above criterion, see section˜8.5. But before coming to the proof, let us show how theorem˜8.1 rules out the exceptional Tannaka group $E_{7}$:
@@
Suppose $G_{X,\omega}^{\ast}\simeq E_{7}$. If $X\subset A$ descends under an isogeny $f\colon A\to B$, then $f(X)\subset B$ is still smooth nondegenerate and irreducible of the same dimension with Tannaka group $E_{7}$ by [JKLM25, rem. 2.6] [Krä22, cor. 1.6(a)]. Hence in what follows we assume that $X$ is nondivisible. Then by [Krä22, cor. 1.10] the representation $V=\omega(\delta_{X})$ of $G_{X,\omega}^{\ast}$ is minuscule and hence isomorphic to the $56$-dimensional irreducible representation of $E_{7}$. By the highest weight theory of the exceptional group $E_{7}$ then
@@
where $W$ is irreducible and $\mathbf{1}$ is the trivial representation of dimension one. In particular, there is an alternating bilinear form on $V$ which is invariant under the action of $E_{7}$. It follows from the parity discussion in [JKLM25, sect. 1.2] that $d$ is odd and, after a suitable translation, that $X$ is symmetric. Then
@@
and $S_{+}(\delta_{X})$ is perverse by the first part of theorem˜8.1 and corresponds to the representation $\omega(S_{+}(\delta_{X}))\simeq W$. The latter is irreducible, so $S_{+}(\delta_{X})$ is a simple perverse sheaf up to negligible direct summands. The second part of theorem˜8.1 then says that $G_{X,\omega}^{\ast}$ is the full symplectic group $\operatorname{Sp}(V,\theta)$, a contradiction. ∎
@@ 8.2. Topological preliminaries
In this section we work over $k=\mathbb{C}$. For the proof of theorem˜8.1 we will detect perverse direct summands of a convolution product from the top cohomology of the fibers of the sum morphism, using the following well-known fact:
@@
The above lemma is about the cohomology with constant coefficients, so we want a criterion to decide when the perverse intersection complex on a singular variety is the constant sheaf. A complex variety $W$ is a rational homology manifold if
@@ 8.3. Kashiwara’s estimate
In this section we continue to work over $k=\mathbb{C}$. We will control negligible summands in convolution products in terms of characteristic varieties. As in [JKLM25, def. 2.2], we attach to any subvariety $Y\subset A$ its projective conormal variety
@@
the union of the characteristic varieties of all its perverse cohomology sheaves. We can then control the characteristic varieties of convolution products K=PPK=P*P by the following incarnation of Kashiwara’s estimate:
@@
Let $P_{1},P_{2}\in\textup{Perv}(A)$. If the Gauss maps $\mathbb{P}\operatorname{Char}(P_{i})\to\mathbb{P}_{A}$ are both finite, then for all $i\in\mathbb{Z}$ and any perverse direct summand $Q\subset^{p}\mathcal{H}^{i}(P_{1}*P_{2})$, the support
$Y\;=\;\operatorname{Supp}(Q)$
is either of general type (in which case $Q$ is not negligible), or $Y=A$.
@@ 8.4. Decomposition of the convolution square
The main point in the proof of theorem˜8.1 will be to deduce from our assumption about the symmetric or alternating square the decomposition of the full tensor square. To explain this we go back to the notation of section˜8.1. Let $T_{-}(\delta_{X})$ denote the direct summand complementary to $T_{+}(\delta_{X})$ in the decomposition of the convolution square $\delta_{X}\ast\delta_{X}$, so that
@@
(2) Part (1) implies that $S_{+}(\delta_{X})$ is perverse. Now, the sum $X\times X\to X+X$ is the composite of the quotient morphism $\pi\colon X\times X\to\operatorname{Sym}^{2}X$ with a unique map $f\colon\operatorname{Sym}^{2}X\to X+X$. Write
@@
and by lemma˜8.4 this sheaf is the perverse intersection complex on $\operatorname{Sym}^{2}X$ up to a shift. The definition of the commutativity constraint for the convolution product in [Wei06, sect. 2.1] moreover shows
@@
The decomposition theorem [BBDG18, th. 6.2.5] together with the fact that $f$ is generically finite implies that $\delta_{X+X}\subset T_{+}(\delta_{X})$. Since by assumption $S_{+}(\delta_{X})$ is a simple perverse sheaf, we see that
@@
Then $f_{*}L_{+}\simeq\mathcal{H}^{-2d}(T_{+}(\delta_{X}))$ has generic rank one on its support $X+X$. Since by base change the generic rank of this direct image is the degree of the generically finite morphism
$f\colon\quad\operatorname{Sym}^{2}X\;\longrightarrow\;X+X,$
it follows that $f$ is birational. This implies that also $\mathcal{H}^{-2d}(T_{-}(\delta_{X}))\simeq f_{*}L_{-}$ has generic rank one on its support $X+X$. The structure of perverse sheaves then forces
@@
where $Q$ is a simple perverse sheaf of generic rank one on its support $X+X$ and $R$ is either zero or a semisimple perverse sheaf with strictly smaller support. Note that $R$ cannot be a skyscraper sheaf: By adjunction [Wei06, cor. 1], any skyscraper summand of $\delta_{X}\ast\delta_{X}$ is supported at a point $a\in A(\mathbb{C})$ with X=aXX=a-X, but since $X$ is nondivisible, there is at most one such skyscraper sheaf and it enters $T_{+}(\delta_{X})$.
@@
by lemma˜8.3. It follows that we have
@@
8.5. Proof of theorem˜8.1
Part (1) of proposition˜8.7 says that $S_{+}(V)$ is a perverse sheaf without negligible direct summands, so it only remains to show that if $S_{+}(\delta_{X})$ is a simple perverse sheaf, then
[…]
But this follows from part (2b) of proposition˜8.7 and from the version of Larsen’s alternative given in [JKLM25, lemma 3.7]. For convenience we recall the simple argument when the subvariety $X\subset A$ is symmetric up to translation, which is the only one needed for the application to $E_{7}$. The symmetry of $X$ implies that the representation $V$ is self-dual and we have an isomorphism $V\otimes V\simeq\operatorname{End}(V)$. Let $\mathfrak{h}\subset\operatorname{End}(V)$ be the Lie algebra of $\operatorname{SO}(V,\theta)$ if $d$ is even and of $\operatorname{Sp}(V,\theta)$ if $d$ is odd. Via the preceding isomorphism we have
@@
The representation $\mathfrak{h}$ of $G_{X,\omega}$ therefore corresponds via Tannaka duality to the perverse sheaf $S_{-}(\delta_{X})$, which is simple by part (2b) of proposition˜8.7. Thus $\mathfrak{h}$ is an irreducible representation of the group $G_{X,\omega}$. On the other hand, the adjoint representation $\mathfrak{g}$ of this group is a subrepresentation of $\mathfrak{h}$, thus $\mathfrak{g}=\mathfrak{h}$. ∎
@@ Appendix A Hodge number estimates
$(-1)^{d-p}\chi(X,\Omega^{p}_{X})\;\geqslant\;\begin{cases}\dim V-d+1&\text{for\penalty 10000\ $p\,\in\,\{0,d\}$},\\[1.00006pt] 2&\text{for\penalty 10000\ $p\,\in\,\{1,d-1\}$},\\[1.00006pt] 1&\text{for\penalty 10000\ $2\leqslant p\leqslant d-2$}.\end{cases}$
@@
When p=d1p=d-1, let $G:=\operatorname{Ker}(\delta_{1})$ so the complex $\mathbb{L}^{p}_{X,V}$ becomes an exact sequence
[…]
@@
$0\longrightarrow\mathcal{O}_{\mathbb{P}}(-d)\otimes\textup{H}^{0}(X,\Omega^{1}_{X})\longrightarrow G\longrightarrow\mathcal{O}_{\mathbb{P}}(j)\longrightarrow 0.$
@@
$\chi(X,\Omega^{d-1}_{X})\;=\;\operatorname{rk}G\;\geqslant 2.$
@@
$\textup{H}^{i}(\mathbb{P},G^{\vee}(j))\;=\;0\quad\text{for all\penalty 10000\ $j\in\mathbb{Z}$ and all\penalty 10000\ $i\in\{1,2,\dots,\dim V-d-1\}$}.$
If $\operatorname{rk}G\leqslant\dim V-d$, then by the Evans-Griffiths theorem [Laz04, ex. 7.3.10] it would follow that $G$ is a direct sum of line bundles. This is seen to be impossible by applying [Lom13, lemma 2.2 (ii)] with $e=d$, $q=\dim V$ and a=ja=-j. Therefore, we have $\operatorname{rk}G\geqslant\dim V-d+1$ and the claim follows since from the above exact sequence $\operatorname{rk}G=\chi(X,\Omega_{X}^{d})$. ∎
@@
$(-1)^{d-p}\chi(X,\Omega^{p}_{X})\;\geqslant\;\begin{cases}g-d+1&\text{for\penalty 10000\ $p\,\in\,\{0,d\}$},\\[1.00006pt] 2&\text{for\penalty 10000\ $p\,\in\,\{1,d-1\}$},\\[1.00006pt] 1&\text{for\penalty 10000\ $2\leqslant p\leqslant d-2$}.\end{cases}$
@@
Since $X\subset A$ has ample normal bundle, we know by [Laz04, prop. 6.3.10(i)] that for nonzero $\omega\in V:=\textup{H}^{0}(A,\Omega^{1}_{A})\subset\textup{H}^{0}(X,\Omega^{1}_{X}),$ the zero locus $Z(\omega)\subset X$ is finite. Hence, $m(V)=d$, and theorem˜A.1 applies. ∎